# Optimal Scaling Needs Optimal Norm

### v1

@@ Abstract
Despite recent progress in optimal hyperparameter transfer under model and dataset scaling, no unifying explanatory principle has been established. Using the Scion optimizer, we discover that joint optimal scaling across model and dataset sizes is governed by a single invariant: the operator norm of the output layer. Across models with up to 1.3B parameters trained on up to 138B tokens, the optimal learning rate/batch size pair $(\eta^{\ast},B^{\ast})$ consistently has the same operator norm value — a phenomenon we term norm transfer. This constant norm condition is necessary but not sufficient: while for each dataset size, multiple $(\eta,B)$ reach the optimal norm, only a unique $(\eta^{\ast},B^{\ast})$ achieves the best loss. As a sufficient condition, we provide the first measurement of $(\eta^{\ast},B^{\ast})$ scaling with dataset size for Scion, and find that the scaling rules are consistent with those of the Adam optimizer. Tuning per-layer-group learning rates also improves model performance, with the output layer being the most sensitive and hidden layers benefiting from lower learning rates. We provide practical insights on norm-guided optimal scaling and release our Distributed Scion (Disco) implementation with logs from over two thousand runs to support research on LLM training dynamics at scale.
@@ 1 Introduction
Recently, an emerging paradigm of norm-based optimization (Bernstein & Newhouse, 2024a; Pethick et al., 2025a) has offered a new lens through which to view training dynamics: it reframes optimization as a process that controls the operator norms of the model’s weight matrices and gradient updates. This perspective enables monitoring of model properties during training, potentially revealing insights deeper than the loss curve alone. This raises a natural question: can the norm-based perspective shed light on how to unify optimal model and dataset size scaling?
@@
  • Unifying invariant for optimal scaling. The operator norm of the output layer $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$ (see Definition 2) for the optimal learning rate $(\eta)$ and batch size $(B)$ configuration has the same value — in other words, is invariant or “transfers” — with both model scaling (in width and depth) and dataset scaling (Fig. 2). We refer to this phenomenon as norm transfer, and it provides a necessary condition for optimality. However, it is not sufficient, as multiple non-optimal $(\eta,B)$ pairs can reach the same optimal norm value (Fig. 3(a)).
@@ 2.1 Background & Terminology
Recently, a fundamental shift in the field of optimal scaling occurred with the work of Yang et al. (2024). It changed the focus from model parametrizations towards the norm perspective by showing that Maximum Update Parametrization ($\mu P$) (Yang et al., 2022) can be derived from a more fundamental principle: enforcing a spectral condition on the model weights and their updates during the training. We briefly explain the idea behind each of them below.
$\mu P$ introduces theoretically grounded scaling rules for hyperparameters as a function of model width in order to ensure “maximal” feature learning in the infinite width limit. This way, the model is guaranteed to learn meaningful features while remaining stable as one scales up its size. As an important by-product, it was found that models with different widths, once parameterized within $\mu P$, all share the same optimal hyperparameters (e.g. learning rate) — therefore allowing for what is known as zero-shot hyperparameter transfer. This property has been extensively used for the past years to ensure optimal model scaling by tuning hyperparameters for a small (proxy) model, and then effortlessly transferring them to a larger one (OpenAI et al., 2024; Gunter et al., 2024; Dey et al., 2024; Meta AI, 2025; Zuo et al., 2025).
@@
where $\text{row}_{i}(.)$ and $\text{col}_{j}(.)$ denote the $i$-th row and $j$-th column of a matrix. In order to control the operator norms, Bernstein & Newhouse (2024a) derived duality maps, i.e. transformation rules of the gradients induced by a given norm. Applying these transformations not only keeps the gradient updates within the required bound (e.g. Eq. 2), but also ensures the steepest descent under the chosen norm (Bernstein & Newhouse, 2024b). For the norms in  eqs. 4, 5 and 6, the corresponding duality maps for the gradient $\bm{G}$} with singular value decomposition (SVD) ${\bm{G}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}$ are:
@@
One prominent example of the norm-based view on model optimization is the Muon optimizer (Jordan et al., 2024), which proved to outperform Adam at scale (Liu et al., 2025; Wang et al., 2025) and showed great performance for models up to 1T parameters (Team et al., 2025). Muon can be viewed as a specific instantiation of Scion: it optimizes hidden layers under $\lVert.\rVert_{\mathrm{RMS}\to\mathrm{RMS}}$ assumption, and uses Adam for the remaining parameters. However, only in the case with no exponential moving average does Adam coincide with the steepest descent in “max-of-max norm” (Bernstein & Newhouse, 2024b). Since this is uncommon in practice, no “natural” norm applies, making Muon hard to analyze through the norm lens. By contrast, Scion naturally incorporates the norm perspective, updating every layer with an assigned, layer-specific norm, making it easy to interpret.
@@ 2.2 Training setup
In all experiments, we use the Llama 3 architecture (Grattafiori et al., 2024) and torchtitan training framework (Liang et al., 2025). Most of the experiments are performed on the small-scale proxy model with a total size of 69M trainable parameters, including input/output embedding layers. For additional ablations in Sec. 3.2, we scale up the model up to $\times 12$ in width (to 1.3B parameters) and up to $\times 32$ in depth (to 168M). Notably, we employ a norm-everywhere approach, inspired by the concept of well-normedness in Large et al. (2024) and the recent line of work (Loshchilov et al., 2025; Kim et al., 2025). Effectively, we ensure that the input $\bm{x}$} to every Linear layer is normalized to ${\bm{x}}||_{\textrm{RMS}}=1$ by a preceding RMSNorm layer without learnable parameters. More details on model configurations are provided in Appendix  A.2 and Appendix A.3.
@@
For pretraining, we use a high-quality partition of the Nemotron-CC dataset (Su et al., 2025), Llama 3 tokenizer (Grattafiori et al., 2024) with a vocabulary size of $128\text{,}256$ (after padding) and a context window of $4096$. All the models are pretrained with the causal language modelling task. Unless stated otherwise, a constant learning rate schedule without warmup and without decay is used. This allows us, for a given set of hyperparameters, to perform a single long run and evaluate progressively larger dataset sizes, rather than conducting several runs for each dataset individually, thereby substantially reducing computational costs (Hu et al., 2024; Hägele et al., 2024).
@@ 2.3 Optimal norm measurement
Our initial intuition was that for a given model and data scale, there is always some optimal norm value, corresponding to some optimal hyperparameter choice. To establish this, we focus on the output layer with the Scion-assigned $\lVert.\rVert_{\mathrm{RMS}\to\infty}$ norm (hereafter referred to as output norm) as being the most natural layer to study. [^6] The output layer is invariant to both width and depth scaling, it is the most sensitive to learning rate tuning (Sec. 3.4), and it can be viewed as a linear classifier on the learned hidden representations. These considerations make us believe that the output layer plays a “representative” role for the entire model, thus making it a distinct layer to analyse. The choice of $\lVert.\rVert_{\mathrm{RMS}\to\infty}$ norm is motivated by Bernstein & Newhouse (2024a) as mapping from a “natural” continuous RMS norm semantics for hidden model representations onto a discrete vocabulary, although we also ablate this in Appendix A.11.2. Since by default we disable momentum and any regularization, we are only left with learning rate ($\eta$) and batch size ($B$) as hyperparameters to tune for optimality.
@@ 3.2 Optimal norm transfer
After analysing learning rate scans across batch sizes, horizons and models of varying width/depth, we visualise results in Fig. 2, with an extended set of plots in Appendix A.7 and Appendix A.13. Each data point corresponds to optimally tuned learning rate $\eta^{*}$ for a given batch size, minimising training loss for that horizon and model. We report our observations below, separately for each direction of scaling.
Figure 2: Training loss vs. output layer norm across batch sizes. (a) Fixed proxy model (69M parameters) while increasing token horizon from $2^{31}$ to $2^{37}$. (b) Fixed token horizon $2^{33}$ while scaling width/depth of the proxy model as indicated in the legend. Each batch size point (increasing from 3232 in $\times 2$ steps, reflected by marker size) has its learning rate optimally tuned. The optimal batch size per horizon/model configuration is indicated by the filled marker. All curves share optimal norm at $7.0\pm 0.2$ across horizons and $7.4\pm 0.2$ across models (grey band).
Data scaling: After profiling across learning rates and plotting optimal norm against batch size, we observe that for a given horizon there is a single optimal batch size with the corresponding optimal output norm $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}=2^{7.0\pm 0.2}$. Intriguingly, this norm value transfers across horizons. We refer to this phenomenon as norm transfer: the optimal $(\eta,B)$ configuration for a given horizon must result in the optimal norm of $\approx 2^{7}$. Also note that the optimal batch size grows with horizon scaling, which we discuss in Sec. 3.3.
@@
Additional ablations: In practice, one is interested in running Scion with non-zero momentum and with a decaying learning rate schedule. We study the impact of these two options in Appendix A.11 and observe that they both show norm transfer. Notably, the addition of momentum largely reduces sensitivity to batch size choice with multiple values resulting in the same optimal norm and loss (Fig. 9). The impact of learning rate decay is also important, as we find it greatly flattens the norm optimum and thus reduces sensitivity to learning rate choice (Fig. 13(b)). Last but not least, in Appendix A.11.2 we find that norm transfer phenomenon occurs not only in $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$ norm, but also in $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\mathrm{RMS}}$ and $\lVert{\bm{W}}_{\mathrm{in}}\rVert_{1\to\mathrm{RMS}}$ for the input embedding.
@@ 3.4 Optimal per-layer-group learning rate
From analysing Fig. 4(a) and additional ones for different batch sizes (Appendix  A.12) we found that the configuration $\eta_{\mathrm{input}}:\eta_{\mathrm{output}}:\eta_{\mathrm{hidden}}=1:1/8:1$ is always among the top 10%. This symmetry simplifies the learning scan and notably contradicts the optimal configurations suggested in Pethick et al. (2025a) and Riabinin et al. (2025). To study dynamics with horizon scaling, we perform the learning rate grid scan same as in Fig. 4(a) but with constraining $\eta_{\mathrm{input}}=\eta_{\mathrm{output}}$  [^9] In terminology of Bernstein & Newhouse (2024a) this corresponds to mass tuning., for the proxy model with $B=512$. Fig. 4(b) illustrates the results, where we see the optimal hidden ratio ($\eta_{\mathrm{input}}/\eta_{\mathrm{hidden}}=1/8$) transfer across horizons, as well as that it brings loss improvement w.r.t. a constant learning rate baseline. Lastly, we note that again, due to the optimizer design, we expect these observations to hold true under model width scaling.
@@ 4 Related Work
Hyperparameters with data scaling  Remains poorly understood theoretically: Smith & Le (2018) showed for SGD how to adjust learning rate and batch size by modelling optimization trajectory as a stochastic differential equation (SDE). Largely, the problem has been approached experimentally by measuring hyperparameter scaling rules as a function of the dataset size (Shen et al., 2024; Hu et al., 2024; Filatov et al., 2025; Bergsma et al., 2025; Li et al., 2025).
$(\eta,B)$ scaling rules  Historically, studies of interaction between learning rate and batch size emerged as an experimental effort to scale batch size without losing performance (Keskar et al., 2017; Goyal et al., 2018; Hilton et al., 2022). Later, a deeper understand has emerged from various theoretical angles: SDE (Malladi et al., 2024), loss curvature (McCandlish et al., 2018), random matrix theory (Granziol et al., 2021).
Norm-based optimization  Starting from the spectral condition (Yang et al., 2024), the approach of transforming gradient updates based on norm assumptions was fully established in Large et al. (2024); Bernstein & Newhouse (2024a), and recently explored in constraining weights themselves (Newhouse et al., 2025). The steepest descent view allowed for connections with manifold learning (Cesista, 2025) and optimizer design (Riabinin et al., 2025). This line of work has led to Muon (Jordan et al., 2024) and Scion (Pethick et al., 2025a; b), along with improvements (Ahn et al., 2025; Amsel et al., 2025), and benchmarks (Wen et al., 2025; Semenov et al., 2025) thereof.

### v2

@@ Abstract
Despite recent progress in optimal hyperparameter transfer under model and dataset scaling, no unifying explanatory principle has been established. For Adam and Scion optimizers, we discover that joint optimal scaling across model and dataset sizes is conditioned on a single invariant: the operator norm of the output layer. Across models with up to 1.3B parameters trained on up to 138B tokens, the optimal learning rate/batch size pair $(\eta^{\ast},B^{\ast})$ consistently has the same operator norm value — a phenomenon we term norm transfer. This constant norm condition is necessary but not sufficient: while for each dataset size, multiple $(\eta,B)$ reach the optimal norm, only a unique $(\eta^{\ast},B^{\ast})$ achieves the best loss. As a sufficient condition, we provide the first measurement of $(\eta^{\ast},B^{\ast})$ scaling with dataset size for Scion, and find that the scaling rules are consistent with those of Adam. Tuning per-layer-group learning rates also improves model performance, with the output layer being the most sensitive and hidden layers benefiting from lower learning rates. We provide practical insights on norm-guided optimal scaling and release our Distributed Scion (Disco) implementation with logs from over two thousand runs to support research on LLM training dynamics at scale.
@@ 1 Introduction
Recently, an emerging paradigm of norm-based optimization (Bernstein and Newhouse, 2024a; Pethick et al., 2025a) has offered a new lens through which to view training dynamics: it reframes optimization as a process that controls the operator norms of the model’s weight matrices and gradient updates. This perspective enables monitoring of model properties during training, potentially revealing insights deeper than the loss curve alone. This raises a natural question: can the norm-based perspective shed light on how to unify optimal model and dataset size scaling?
@@
  • Unifying invariant for optimal scaling. The operator norm of the output layer $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}$ (see Definition 2) for the optimal learning rate $(\eta)$ and batch size $(B)$ configuration has the same value — in other words, is invariant or “transfers” — with both model scaling (in width and depth) and dataset scaling (Fig. 2), as observed for both Scion and Adam optimizers (Appendix A.12). We refer to this phenomenon as norm transfer, and it provides a necessary condition for optimality. However, it is not sufficient, as multiple non-optimal $(\eta,B)$ pairs can reach the same optimal norm value (Fig. 3(a)).
@@ 2.1 Background & Terminology
Recently, a fundamental shift in the field of optimal scaling occurred with the work of Yang et al. (2024). It changed the focus from model parametrizations towards the norm perspective by showing that Maximum Update Parametrization ($\mu P$) (Yang et al., 2022) can be derived from a more fundamental principle: enforcing a spectral condition on the model weights and their updates during the training. We briefly explain the idea behind these concepts below.
$\mu P$ introduces theoretically grounded scaling rules for hyperparameters as a function of model width in order to ensure “maximal” feature learning in the infinite width limit. This way, the model is guaranteed to learn meaningful features while remaining stable as one scales up its size. As an important by-product, it was found that models with different widths, once parameterized within $\mu P$, all share the same optimal hyperparameters (e.g. learning rate) — therefore allowing for what is known as zero-shot hyperparameter transfer. This property has been extensively used for the past years to ensure optimal model scaling by tuning hyperparameters for a small (proxy) model, and then effortlessly transferring them to a larger one (Gunter et al., 2024; Dey et al., 2024; Meta AI, 2025; Zuo et al., 2025).
@@
where $\text{row}_{i}(.)$ and $\text{col}_{j}(.)$ denote the $i$-th row and $j$-th column of a matrix. In order to control the operator norms, Bernstein and Newhouse (2024a) derived duality maps, i.e. transformation rules of the gradients induced by a given norm. Applying these transformations not only keeps the gradient updates within the required bound (e.g. Eq. 2), but also ensures the steepest descent under the chosen norm (Bernstein and Newhouse, 2024b). For the norms in  eqs. 4, 5 and 6, the corresponding duality maps for the gradient $\bm{G}$} with singular value decomposition (SVD) ${\bm{G}}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{\top}$ are:
@@
One prominent example of the norm-based view on model optimization is the Muon optimizer (Jordan et al., 2024), which proved to outperform Adam at scale (Liu et al., 2025; Wang et al., 2025) and showed great performance for models up to 1T parameters (Team et al., 2025). Muon can be viewed as a specific instantiation of Scion: it optimizes hidden layers under $\lVert.\rVert_{\mathrm{RMS}\to\mathrm{RMS}}$ assumption, and uses Adam for the remaining parameters. However, only in the case with no exponential moving average does Adam coincide with the steepest descent in “max-of-max norm” (Bernstein and Newhouse, 2024b). Since this is uncommon in practice, no “natural” norm applies, making Muon hard to analyze through the norm lens. By contrast, Scion naturally incorporates the norm perspective, updating every layer with an assigned, layer-specific norm, making it easy to interpret.
@@ 2.2 Training setup
In all experiments, we use the Llama 3 architecture (Grattafiori et al., 2024) and torchtitan training framework (Liang et al., 2025). Most of the experiments are performed on the model with a total size of 69M trainable parameters (including input/output embedding layers), hereafter referred to as proxy model. For additional ablations in Sec. 3.2, we scale up the model up to $\times 12$ in width (to 1.3B parameters) and up to $\times 32$ in depth (to 168M). Notably, we employ a norm-everywhere approach, inspired by the concept of well-normedness in Large et al. (2024) and the recent line of work (Loshchilov et al., 2025; Kim et al., 2025). Effectively, we ensure that the input $\bm{x}$} to every Linear layer is normalized to ${\bm{x}}||_{\textrm{RMS}}=1$ by a preceding RMSNorm layer without learnable parameters. More details on model configurations are provided in Appendix  A.2 and Appendix A.3.
@@
For pretraining, we use a high-quality partition of the Nemotron-CC dataset (Su et al., 2025), Llama 3 tokenizer (Grattafiori et al., 2024) with a vocabulary size of $128\text{,}256$ (after padding) and a context window of $4096$. All the models are pretrained with the causal language modelling task. Unless stated otherwise, a constant learning rate schedule without warmup and without decay is used. This allows us, for a given set of hyperparameters, to perform a single long run and evaluate progressively larger dataset sizes, rather than conducting several runs for each dataset individually, thereby substantially reducing computational costs (Hu et al., 2024; hägele2024scalinglawscomputeoptimaltraining).
@@ 2.3 Optimal norm measurement
Our initial intuition was that for a given model and data scale, there is always some optimal norm value, corresponding to some optimal hyperparameter choice. To establish this, we focus on the output layer with the Scion-assigned $\lVert.\rVert_{\mathrm{RMS}\to\infty}$ norm (hereafter referred to as output norm) as being the most natural layer to study. [^6] The output layer is invariant to both width and depth scaling, it is the most sensitive to learning rate tuning (Sec. 3.4), and it can be viewed as a linear classifier on the learned hidden representations. These considerations make us believe that the output layer plays a “representative” role for the entire model, thus making it a distinct layer to analyse. The choice of $\lVert.\rVert_{\mathrm{RMS}\to\infty}$ norm is motivated by Bernstein and Newhouse (2024a) as mapping from a “natural” continuous RMS norm semantics for hidden model representations onto a discrete vocabulary, although we also ablate this in Appendix A.15.2. Since by default we disable momentum and any regularization, we are only left with learning rate ($\eta$) and batch size ($B$) as hyperparameters to tune for optimality.
@@ 3.2 Optimal norm transfer
After analysing learning rate scans across batch sizes, horizons and models of varying width/depth, we visualise results in Fig. 2, with an extended set of plots in Appendix A.7 and Appendix A.17. Each data point corresponds to optimally tuned learning rate $\eta^{*}$ for a given batch size, minimising training loss for that horizon and model. We report our observations below, separately for each direction of scaling, as well as additional ablations.
Figure 2: Training loss vs. output layer norm across batch sizes. (a) Fixed proxy model (69M parameters) while increasing token horizon from $2^{31}$ to $2^{37}$. (b) Fixed token horizon $2^{33}$ while scaling width/depth of the proxy model as indicated in the legend. Each batch size point (increasing from $32$ in $\times 2$ steps, reflected by marker size) has its learning rate optimally tuned. The optimal batch size per horizon/model configuration is indicated by the filled marker. All curves share optimal norm at $7.0\pm 0.2$ across horizons and $7.4\pm 0.2$ across models (grey band).
Data scaling: After profiling across learning rates and plotting optimal norm against batch size, we observe that for a given horizon there is a single optimal batch size with the corresponding optimal output norm $\lVert{\bm{W}}_{\mathrm{out}}\rVert_{\mathrm{RMS}\to\infty}=2^{7.0\pm 0.2}$. Intriguingly, this norm value transfers across horizons. We refer to this phenomenon as norm transfer: the optimal $(\eta,B)$ configuration for a given horizon must result in the optimal norm of $\approx 2^{7}$. Also note that the optimal batch size grows with horizon scaling, which we discuss in Sec. 3.3. Interestingly, we observe the same norm transfer behavior when switching to a different dataset (Appendix A.11): specifically, Thai and Russian language partitions of Fineweb-2 (Penedo et al., 2025).
@@
Momentum & learning rate decay: In practice, one is more interested in Scion with non-zero momentum and with a decaying learning rate schedule as resulting in better performance. We study the impact of these two options in Appendix A.15 and observe that they both show norm transfer. Notably, the addition of momentum largely reduces sensitivity to batch size choice with multiple values resulting in the same optimal norm and loss (Fig. 14). The same is applicable to learning rate decay, which reduces sensitivity to learning rate choice (Fig. 18(b)).
Adam optimizer: As the optimizer commonly used in practice, we study its data scaling for the proxy model with two configurations: with momenta ($\beta_{1}=0.9,\beta_{2}=0.95$) and without ($\beta_{1}=\beta_{2}=0$), where the latter case corresponds to a sign gradient descent. Intriguingly, for both we observe a clear norm transfer (Appendix A.12): the case without momentum exhibits the same optimal norm value ($\approx 2^{7}$) as the analogous without-momentum Scion, while the case with momentum has a noticeably higher optimal norm ($\approx 2^{11}$). We believe that this shared norm transfer pattern further supports a common norm-based view on optimization by Bernstein and Newhouse (2024b), and therefore makes our observations for Scion also transferable to Adam.
Normalization layers: Since we are interested in the norm structure of model scaling, our choice of norm-everywhere strategy may play a key role in the observed phenomena. We ablate various ways to place normalization layers (RMSNorm without trainable parameters) within the architecture in Appendix A.14. For the proxy model, fixed data horizon of 43B tokens, and fixed batch size $B=256$ sequences we perform a learning rate scan while removing specific normalization layers (see Appendix A.14 for a detailed description). In Fig. 13(b) for the case of the Scion optimizer with momentum $\alpha=0.1$ we observe that removing normalization in residual connections results in significant training divergences. The setup with QK-norm + residuals + output layer normalization results in the same learning rate profile as our default norm-everywhere, indicating redundancy of MLP- and VO-normalization. Furthermore, residuals + output configuration results in slightly better performance, albeit with higher learning rate sensitivity. Likewise, addition of QK-norm largely reduces learning rate sensitivity, as also shown by Wortsman et al. (2023), thus making the training less sensitive to learning rate choice.
Finally, we selected the residuals-only configuration as the best trade-off between the minimalist usage of normalization layers, learning rate sensitivity, and model performance, and studied its norm scaling properties. Fig. 12 shows that norm transfer is present also in this scenario, interestingly with significantly lower optimal norm comparing to the norm-everywhere setting ($\approx 2^{3}$, one order of magnitude lower). One can also notice from Fig. 13(b) (see norm values in the legend) that it is the removal of the output layer norm that induces this reduction. This finding poses an interesting question: can it be in any way advantageous to prefer the model with lower weight norms? Intuitively, we would answer positively, referring to discussions in Newhouse et al. (2025), but leave detailed studies for future work.
@@ 3.4 Optimal per-layer-group learning rate
From analysing Fig. 4(a) and additional ones for different batch sizes (Appendix  A.16) we found that the configuration $\eta_{\mathrm{input}}:\eta_{\mathrm{output}}:\eta_{\mathrm{hidden}}=1:1/8:1$ is always among the top 10%. This symmetry simplifies the learning scan and notably contradicts the optimal configurations suggested in Pethick et al. (2025a) and Riabinin et al. (2025). To study dynamics with horizon scaling, we perform the learning rate grid scan same as in Fig. 4(a) but with constraining $\eta_{\mathrm{input}}=\eta_{\mathrm{output}}$  [^9] In terminology of Bernstein and Newhouse (2024a) this corresponds to mass tuning., for the proxy model with $B=512$. Fig. 4(b) illustrates the results, where we see the optimal hidden ratio ($\eta_{\mathrm{input}}/\eta_{\mathrm{hidden}}=1/8$) transfer across horizons, as well as that it brings loss improvement w.r.t. a constant learning rate baseline. Lastly, we note that again, due to the optimizer design, we expect these observations to hold true under model width scaling.
@@ 4 Related Work
Hyperparameters with data scaling  Remains poorly understood theoretically: Smith and Le (2018) showed for SGD how to adjust learning rate and batch size by modelling optimization trajectory as a stochastic differential equation (SDE). Largely, the problem has been approached experimentally by measuring hyperparameter scaling rules as a function of the dataset size (Shen et al., 2024; Hu et al., 2024; Filatov et al., 2025; Bergsma et al., 2025; Li et al., 2025).
$(\eta,B)$ scaling rules  Historically, studies of interaction between learning rate and batch size emerged as an experimental effort to scale batch size without losing performance (Keskar et al., 2017; Goyal et al., 2018; Hilton et al., 2022). Later, a deeper understanding has been built from various theoretical angles: SDE (Malladi et al., 2024; Compagnoni et al., 2024), loss curvature (McCandlish et al., 2018), random matrix theory (Granziol et al., 2021).
Norm-based optimization  Starting from the spectral condition (Yang et al., 2024), the approach of transforming gradient updates based on norm assumptions was fully established in Large et al. (2024); Bernstein and Newhouse (2024a), and recently explored in constraining weights themselves (Newhouse et al., 2025). The steepest descent view allowed for connections with manifold learning (Cesista, 2025) and optimizer design (Riabinin et al., 2025). This line of work has led to Muon (Jordan et al., 2024) and Scion (Pethick et al., 2025a; b), along with improvements (Ahn et al., 2025; Amsel et al., 2025), and benchmarks (Wen et al., 2025; Semenov et al., 2025) thereof.